7. A nearsighted man uses his ideally prescribed eyeglasses that have a refractive power of −2.50 diopters. He would like a prescription for contact lenses. What is the correct contact lens prescription if he normally wears his eyeglasses a distance of 2.8 cm from his eyes?
Answer #1

Assume the glasses are 2.8 cm from the person eyes and the person near-point distance is 25 cm when wearing the glasses.

The image distance for distant viewing which is given as -

using a len's formula, we have

1 / d0 + 1 / di = 1 / f

1 / di = (-2.5 m) + (0 m)

di = - (1 / 2.5) m

di = - 0.4 m

Add the eye-to-lens distance

Far point = - di + (0.028 m)

Far point = [(0.4 m) + (0.028 m)]

Far point = 0.428 m

With the glasses, the person can see clearly for all distances greater than 25 cm.